OpenStudy (anonymous):

Please help me!!! I need it! Sally has only nickels and dimes in her money box. She knows that she has less than $20 in the box. Let x represent the number of nickels in the box and y represent the number of dimes in the box. Which of the following statements best describes the steps to graph the solution to the inequality in x and y?

OpenStudy (anonymous):

Draw a dashed line to represent the graph of 5x + 10y = 2000 and shade the portion above the line for positive values of x and y. Draw a dashed line to represent the graph of 10x + 5y = 2000 and shade the portion above the line for positive values of x and y. Draw a dashed line to represent the graph of 5x + 10y = 2000 and shade the portion below the line for positive values of x and y. Draw a dashed line to represent the graph of 10x – 5y = 2000 and shade the portion below the line for positive values of x and y.

OpenStudy (anonymous):

@Whitemonsterbunny17 This is way to confusing! Can you help? :)

whitemonsterbunny17 (whitemonsterbunny17):

Which part is confusing? the number part or where the shading goes?

whitemonsterbunny17 (whitemonsterbunny17):

or both? lol

OpenStudy (anonymous):

The shading part sorry. :)

whitemonsterbunny17 (whitemonsterbunny17):

ohhh okay.. we'll narrow it down by deciding which one you think it is with the number part?

OpenStudy (anonymous):

I think it would be 5x+10y=2000

OpenStudy (anonymous):

Would it'd be C?

whitemonsterbunny17 (whitemonsterbunny17):

Yeah, because you want it to be less than $20.. so for less than you would shade below the line. If it was greater than, you would shade above the line. (:

OpenStudy (anonymous):

Thanks soo much! I appreciate it. ♥☻

whitemonsterbunny17 (whitemonsterbunny17):

You're welcome! =^.^=

OpenStudy (anonymous):

:)